OpenStudy (anonymous):

What is the sum of the arithmetic sequence 140, 134, 128…, if there are 33 terms?

OpenStudy (anonymous):

a=140, d=-6, n=33. sum=n/2(2a+(n-1)d) =33/2(280+32x(-6)) =33/2(88) =1452